OpenStudy (anonymous):

A flat circular plate of copper has a radius of 0.425 m and a mass of 65 kg. What is the thickness of the plate? Answer in units of m. I would like you to guide me to the answer. I switched the numbers up so they are not the exact as on my paper (: Thanks.

OpenStudy (anonymous):

first you need the density of copper and then multiply it by the area of the plate

OpenStudy (anonymous):

Density of copper is 8.9. Check.

OpenStudy (anonymous):

the area of the plate is (pi)r^2 so (.425m)^2 (pi)

OpenStudy (anonymous):

.180625pi meters squared=area

OpenStudy (anonymous):

density is 8.9 kg/m2? or g/m2?

OpenStudy (anonymous):

No wait! I take that back o: I am confusing mystel. 8.9=g/cm^3 And 8900=kg/m^3

OpenStudy (anonymous):

The thickness times area is volume. times the density is the mass. you are given mass so the thickness is equal to the density divided by area

OpenStudy (anonymous):